3. Two turtles are racing. The length of time that turtle A takes is expo 2 nentially distributed with mean 5 minutes. The length that turtle B take is also exponentially distributed but with mean 7 minutes. Assume tha their times are independent. (a) What is the probability that A wins? (b) What is the probability that the winner takes longer than 6 minutes (c) What is the expected time of the winner? (d) What is the probability of a tie?
Answer #1

Let X be random variable denoting length of time that turtle A takes Y be random variable denoting length of time that turtle

(c) The cdf of = min(x, y is Fu(u) = P(min{X,Y} < u) = 1-P(X > u, Y > u) exp 12u The pdf of U is dFv(u) 12 0 otherwise Hence